3x+3y+=2
2
3X-4Y=-16
im stuck i just stared algebra and i suck!

Answer by bacos(9)

Simplifying
3X + -4Y = -16
Solving
3X + -4Y = -16
Solving for variable 'X'.
Move all terms containing X to the left, all other terms to the right.
Add '4Y' to each side of the equation.
3X + -4Y + 4Y = -16 + 4Y
Combine like terms: -4Y + 4Y = 0
3X + 0 = -16 + 4Y
3X = -16 + 4Y
Divide each side by '3'.
X = -5.333333333 + 1.333333333Y
Simplifying
X = -5.333333333 + 1.333333333Y